find the gradients
2x+3y=7
4y-5x=2
i need it hurry our great expert

Answer by macston(5194)

.
Quickest way is to convert to slope intercept form:
y=mx+b, where m is the slope (gradient)
.
Equation 1:
2x+3y=7 Subtract 2x from each side.
3y=-2x+7 Divide each side by 3.
y=(-2x+7)/3
y=(-2/3)x+(7/3)
The slope (or gradient,m) is -2/3.
.
Equation 2:
4y-5x=2 Add 5x to each side.
4y=5x+2 Divide each side by 4.
y=(5x+2)/4
y=(5/4)x+1/2
The slope is 5/4